OpenStudy (anonymous):

Solve the following quadratic equation using the quadratic formula and then choose the correct solution set. 5x^2 + 7x = 6 {-2, 3/5} ?

OpenStudy (anonymous):

{-2, 3/5} is correct.

OpenStudy (anonymous):

could you explain the steps?

OpenStudy (anonymous):

OK, wait

OpenStudy (anonymous):

Equate to zero \[=> 5x^2 + 7x = 6\] \[=> 5x^2 + 7x - 6= 0\] Quad. Equation - \[(-b ± \sqrt{b^2 -4ac})/2a\] Plugin the numbers - \[(-7 ± \sqrt{7^2 -(4*5*-6)})/2*5\] Solve - \[(-7 ± \sqrt{49 -(-120)})/10\] \[(-7 ± \sqrt{169})/10\] \[(-7 ± 13)/10\] Root 1 = \[(-7 + 13)/10\] = 3/5 Root 2 = \[(-7 - 13)/10\] = -20/10 = -2 *Sorry for weird formatting as i'm new to this type of typesetting.

OpenStudy (anonymous):

okay, thanks.

OpenStudy (anonymous):

gotit.